OpenStudy (anonymous):

solve each system of equations by substitution x-y=1 x+y=-1

OpenStudy (anonymous):

1. Set one equation to y 2. Substitute the y in the other equation. 3. Solve

OpenStudy (anonymous):

x-y=1 ---> y=x-1 x+ (x-1) =-1 2x-1=-1 2x=0 x=0